# RoboRefer: Towards Spatial Referring with Reasoning in Vision-Language Models for Robotics

### v3

@@ E.3 Spatial Referring Benchmarks
We evaluate three recent robotic referring benchmarks—RoboRefIt [137], Where2Place [5], and RoboSpatial [2], all limited to $2$ reasoning steps. Specifically, RoboRefIt concentrates on object location referring by leveraging object attributes and spatial relations with anchor objects. Where2Place further explores how to place objects relative to anchor objects in the camera’s coordinate frame. RoboSpatial builds on this idea, considering the same form of placement relative to anchor objects, but many samples in benchmarks consider the object-centric coordinate frame. We also evaluate more complex multi-step spatial referring on RefSpatial-Bench, a challenging benchmark based on real-world cluttered scenes, as introduced in Appx. C. For all these benchmarks, we use the same evaluation protocol: we first compute the proportion of predicted points that fall within the ground truth mask for each sample, and then average the results across all samples to obtain the success rate.
@@ E.4 Simulation Evaluation
We use the same evaluation protocol of Open6DOR V2 introduced in SoFar [7], following the official repositoryhttps://github.com/Zhangwenyao1/Open6DOR_V2_Execution. Specifically, we only test the position track as this work focuses on location and placement via spatial referring rather than executing 6DOF manipulation tasks. Notably, we find that our model achieves nearly 100% success in the perception stage (i.e., determining location and placement), with failures primarily attributed to motion planning errors such as IK failures or collision-prone trajectories. We show more demonstrations of simulation evaluation in Appx. F.
@@ E.5.1 UR5 Manipulation
We show two demos for UR5 https://www.universal-robots.com/products/ur5e/ Manipulation: human disturbance and voice interruption. In the human disturbance case, RoboRefer runs at 2.5Hz. Significant shifts in predicted 2D coordinates trigger motion interruption and re-planning. In the voice interruption case, incoming speech commands are continuously monitored. Upon detection, the current task is halted. We use the Whisper [155] ASR model to transcribe speech, which RoboRefer processes into new 2D coordinates for task redirection.
For grasping, the 2D coordinates are fed into SAM2 [77] to generate a segmentation mask, which filters the target object’s point cloud from the scene captured by a third-person Intel RealSense L515 https://www.intelrealsense.com/lidar-camera-l515/ depth camera. The extracted point cloud is input to AnyGrasp [156] to predict a grasp pose in the camera coordinate frame. Using an eye-to-hand calibration method, the grasp pose is transformed into the UR5 robot’s base frame for execution.
@@ E.5.2 G1 Humanoid Mobile Manipulation
For grasping, we employ a head-mounted Intel RealSense D435https://www.intelrealsense.com/depth-camera-d435/ on the Unitree G1 humanoid to capture RGB-D images, which are processed by RoboRefer to extract 2D target coordinates. These coordinates guide SAM2 [77] to generate a segmentation mask, which filters the third-person D435 point cloud to isolate the target object. The filtered point cloud is then passed to AnyGrasp [156] to predict a grasp pose in the third-person frame, which is transformed to the robot’s base frame using known camera-to-robot calibration.
@@
Method Input CV-Bench [15] $\text{BLINK}_{val}$ [16]
@@ Appendix G More Discussion on Limitations and Future Work
Our model struggles with these cases because RefSpatial lacks multi-step referring data that embeds human priors and intent understanding. As RefSpatial is procedurally generated, incorporating such characteristics at scale remains challenging. Future work may explore procedural synthesis of intent-aware data or improve model performance via co-training with datasets supporting intent comprehension, such as PixMo-Points [10].
@@ Appendix I Licenses
(1) 2D web image data: OpenImages [132] is released under Apache License 2.0.
(2) 3D embodied video data: CA-1M [133] is released under CC-by-NC-ND.
(3) Procedural scene generation: Infinigen [134] is released under BSD 3-Clause License.
(4) 3D digital assets: objaverse-xl [135] is released under Apache License 2.0.
@@
Figure 40: Visualization of RoboRefer’s prediction (blue point) on Open6DOR V2 [7] benchmark.

### v4

@@ E.3 Spatial Referring Benchmarks
We evaluate three recent robotic referring benchmarks—RoboRefIt [lu2023vl], Where2Place [yuan2024robopoint], and RoboSpatial [song2024robospatial], all limited to $2$ reasoning steps. Specifically, RoboRefIt concentrates on object location referring by leveraging object attributes and spatial relations with anchor objects. Where2Place further explores how to place objects relative to anchor objects in the camera’s coordinate frame. RoboSpatial builds on this idea, considering the same form of placement relative to anchor objects, but many samples in benchmarks consider the object-centric coordinate frame. We also evaluate more complex multi-step spatial referring on RefSpatial-Bench, a challenging benchmark based on real-world cluttered scenes, as introduced in Appx. C. For all these benchmarks, we use the same evaluation protocol: we first compute the proportion of predicted points that fall within the ground truth mask for each sample, and then average the results across all samples to obtain the success rate.
@@ E.4 Simulation Evaluation
We use the same evaluation protocol of Open6DOR V2 introduced in SoFar [qi2025sofar], following the official repositoryhttps://github.com/Zhangwenyao1/Open6DOR_V2_Execution. Specifically, we only test the position track as this work focuses on location and placement via spatial referring rather than executing 6DOF manipulation tasks. Notably, we find that our model achieves nearly 100% success in the perception stage (i.e., determining location and placement), with failures primarily attributed to motion planning errors such as IK failures or collision-prone trajectories. We show more demonstrations of simulation evaluation in Appx. F.
@@ E.5.1 UR5 Manipulation
We show two demos for UR5 https://www.universal-robots.com/products/ur5e/ Manipulation: human disturbance and voice interruption. In the human disturbance case, RoboRefer runs at 2.5Hz. Significant shifts in predicted 2D coordinates trigger motion interruption and re-planning. In the voice interruption case, incoming speech commands are continuously monitored. Upon detection, the current task is halted. We use the Whisper [radford2023robust] ASR model to transcribe speech, which RoboRefer processes into new 2D coordinates for task redirection.
For grasping, the 2D coordinates are fed into SAM2 [ravi2024sam] to generate a segmentation mask, which filters the target object’s point cloud from the scene captured by a third-person Intel RealSense L515 https://www.intelrealsense.com/lidar-camera-l515/ depth camera. The extracted point cloud is input to AnyGrasp [fang2023anygrasp] to predict a grasp pose in the camera coordinate frame. Using an eye-to-hand calibration method, the grasp pose is transformed into the UR5 robot’s base frame for execution.
@@ E.5.2 G1 Humanoid Mobile Manipulation
For grasping, we employ a head-mounted Intel RealSense D435https://www.intelrealsense.com/depth-camera-d435/ on the Unitree G1 humanoid to capture RGB-D images, which are processed by RoboRefer to extract 2D target coordinates. These coordinates guide SAM2 [ravi2024sam] to generate a segmentation mask, which filters the third-person D435 point cloud to isolate the target object. The filtered point cloud is then passed to AnyGrasp [fang2023anygrasp] to predict a grasp pose in the third-person frame, which is transformed to the robot’s base frame using known camera-to-robot calibration.
@@
Method Input CV-Bench [tong2024cambrian] $\text{BLINK}_{val}$ [fu2024blink]
@@ Appendix G More Discussion on Limitations and Future Work
Our model struggles with these cases because RefSpatial lacks multi-step referring data that embeds human priors and intent understanding. As RefSpatial is procedurally generated, incorporating such characteristics at scale remains challenging. Future work may explore procedural synthesis of intent-aware data or improve model performance via co-training with datasets supporting intent comprehension, such as PixMo-Points [deitke2024molmo].
@@ Appendix I Licenses
(1) 2D web image data: OpenImages [kuznetsova2020open] is released under Apache License 2.0.
(2) 3D embodied video data: CA-1M [lazarow2024cubify] is released under CC-by-NC-ND.
(3) Procedural scene generation: Infinigen [raistrick2024infinigen] is released under BSD 3-Clause License.
(4) 3D digital assets: objaverse-xl [deitke2023objaverse] is released under Apache License 2.0.
@@
Figure 40: Visualization of RoboRefer’s prediction (blue point) on Open6DOR V2 [qi2025sofar] benchmark.